# Bounds on the number of squares in recurrence sequences

### v1

@@ 3.1. Representation Proposition
From Lemma 3.3(a), $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of $4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)$. We put $r=r_{1}/g_{1}$, for either choice of $r_{1}$, $s=s_{1}^{\prime}/g_{1}$ and $f=4b^{2}r_{1}/g_{1}^{2}$. By the definition of $g_{1}^{2}$, we see that $f,r,s\in\mathbb{Z}$.
Lastly, we choose $r_{1}$. Let $d^{\prime}$ be the product of all primes, $p|\operatorname{sf}\left(N_{\alpha}\right)$, such that the minimum in Lemma 3.3(b) is attained for $r_{1}^{\prime}/{g_{1}^{\prime}}^{2}$ and $d^{\prime\prime}$ be the product of the remaining primes, $p|\operatorname{sf}\left(N_{\alpha}\right)$ (so the minimum in Lemma 3.3(b) is attained for $r_{1}^{\prime\prime}/{g_{1}^{\prime\prime}}^{2}$, but not for $r_{1}^{\prime}/{g_{1}^{\prime}}^{2}$ for such primes). If $d^{\prime}>d^{\prime\prime}$, then we let $r_{1}=r_{1}^{\prime}$ and let $r_{1}=r_{1}^{\prime\prime}$ otherwise. This choice ensures that $f^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)/\max\left(d^{\prime},d^{\prime\prime}\right)$ satisfies the desired conditions.
(b) If $f$ is odd, then there is nothing to prove, as $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$ follows from $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. So we need only consider when $f$ is even.
@@
We can write $(r+si)/(1+i)=((r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
$-\left(\frac{2b^{2}r}{g_{1}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
Since $f=4b^{2}r/g_{1}$ is even, we know that $2b^{2}r/g_{1}\in\mathbb{Z}$.
Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. Hence $2b^{2}r/g_{1}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)$.
@@
(b-ii-1) Suppose first that $v_{2}\left(r_{1}\right)$ is even. In this case, $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$. If we have equality, then $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and so $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required. So we consider $v_{2}\left(4b^{2}\right)>v_{2}\left(r_{1}\right)$.
Since both of these quantities are even, we must have $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)+2$. That is, $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$.
Since $s$ is even, we have $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)>2v_{2}\left(g_{1}\right)=2v_{2}\left(r_{1}^{\prime}\right)$. So $v_{2}\left(r_{1}^{\prime\prime}\right)>v_{2}\left(r_{1}^{\prime}\right)$. As we stated near the start of this case ((b-ii)), since $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$, we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
We saw in the proof of Lemma 3.2 that $v_{2}(y)\geq\min\left(v_{2}(b),v_{2}\left(r_{1}^{\prime}\right)/2\right)$. Since $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$, it follows that $v_{2}(y)\geq v_{2}\left(r_{1}^{\prime}\right)/2$. Hence $v_{2}(2by)\geq 1+v_{2}\left(r_{1}\right)/2+v_{2}\left(r_{1}\right)/2=v_{2}\left(r_{1}\right)+1$. Hence $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)\geq v_{2}\left(r_{1}\right)+1$, which implies that $v_{2}\left(r_{1}\right)>v_{2}\left(r_{1}\right)$. Therefore, the case when $v_{2}\left(r_{1}\right)$ is even and $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$ never occurs and we find that in case (b-ii-1), $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
(b-ii-2) Now suppose that $v_{2}\left(r_{1}\right)$ is odd. Here $v_{2}\left(2b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$.
@@
Since $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)=v_{2}\left(g_{1}^{2}\right)<2v_{2}\left(r_{1}\right)$. Hence $v_{2}\left(r_{1}^{\prime\prime}\right)\neq v_{2}\left(r_{1}\right)$, so by the comment at the start of case (b-ii), we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
(b-iii-1) If $v_{2}\left(r_{1}\right)$ is even, then $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}\right)$. Hence $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
@@
Since $s$ is even, we have $2\leq v_{2}\left(s^{2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)-v_{2}\left(2b^{2}r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime\prime}\right)-v_{2}\left(2b^{2}\right)$. So $v_{2}\left(r_{1}^{\prime\prime}\right)\geq 2v_{2}(b)+3$.
@@
From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4bN_{\varepsilon}$. So $atu+du^{2}$ must be even. I.e., $atu$ and $du^{2}$ have the same parity.
If $du^{2}$ is odd, then $a$ and $t$ are also odd. Since $4N_{\varepsilon}$ is divisible by $4$, it must be the case that $d\equiv 1\pmod{4}$. If $d\equiv 1\pmod{8}$, then $N_{\varepsilon}$ is even. Furthermore, since $abd$ is odd, $N_{\alpha}$ is even, so $v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)\geq 1$ and the desired conclusion holds.
@@
We now consider $du^{2}$ even. Here we will prove that both $uN_{\alpha}$ and $N_{\varepsilon}$ are even. This suffices to prove part (b) in this case.
@@
Since $4N_{\varepsilon}=t^{2}-du^{2}\in\mathbb{Z}$ and $du^{2}$, we have $t$ is even and hence $4|\left(du^{2}\right)$.
If $u$ is odd, then $4|d$. Also $N_{\alpha}$ is even, so $a$ is even. As a result we find that $v_{2}(2atu),v_{2}\left(2b^{2}du^{2}\right)\geq 3$. So under our assumption that $N_{\varepsilon}$ is odd, we have $v_{2}\left(4y^{2}\right)=2$. I.e., $y$ is odd.
@@
In both cases, we find that $y$ is even. This contradicts the assumption that $N_{\varepsilon}$ is odd. Hence we know that in the case when $du^{2}$ is even, $N_{\varepsilon}$ is even and the assertion in part (b) of the lemma holds.
@@
(c-i) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=p$, for an odd prime, $p$. Then we have $f^{\prime}=1$ in part (a) and so $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$. Hence the result holds if $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
From Lemma 3.4, $r$ is even unless $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$. So, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$. With $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(N_{\varepsilon}\right)\right)$.
(c-ii) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=2p$, where $p$ is an odd prime. Here the worst case is when the minimum in (3.7) occurs for one of $r_{1}^{\prime}/g_{1}^{\prime 2}$ and $r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}$ for $2$ and occurs for the other for $p$. In this case, we let $r$ be the one of $r_{1}^{\prime}$ or $r_{1}^{\prime\prime}$ for which the minimum in (3.7) for $p$ occurs. Thus $v_{p}\left(r/g_{1}^{2}\right)\leq v_{p}\left(\operatorname{rad}\left(N_{\varepsilon}\right)\right)$ for all odd primes, $p$. From Lemma 3.4, we see that $r$ is even. Hence, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$. So $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(N_{\varepsilon}\right)\right)$. ∎

### v2

@@ 3.1. Representation Proposition
From Lemma 3.3(a), $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of
$4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=4b^{2}\operatorname{rad}\left(\operatorname{sf}\left(N_{\alpha}\right)\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right).$
We put $r=r_{1}/g_{1}$, for either choice of $r_{1}$, $s=s_{1}^{\prime}/g_{1}$ and $f=4b^{2}r_{1}/g_{1}^{2}$. By the definition of $g_{1}^{2}$, we see that $f,r,s\in\mathbb{Z}$.
Lastly, we choose which of the two possible values of $r_{1}$ to use. As in the statement of Lemma 3.3, we let $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$ be the two possibilities. We will use Lemma 3.3(b). We divide the prime divisors of $\operatorname{sf}\left(N_{\alpha}\right)$ into three disjoint sets:
$I^{\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\},$
$I^{\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)<v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\right\},$
$I^{\prime\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\}.$
With $d^{\prime}=\prod_{p\in I^{\prime}}p$, $d^{\prime\prime}=\prod_{p\in I^{\prime\prime}}p$ and $d^{\prime\prime\prime}=\prod_{p\in I^{\prime\prime\prime}}p$, we have $d^{\prime}d^{\prime\prime}d^{\prime\prime\prime}=\operatorname{sf}\left(N_{\alpha}\right)$.
We put $r_{1}=r_{1}^{\prime}$ if $d^{\prime}>d^{\prime\prime}$ and $r_{1}=r_{1}^{\prime\prime}$ otherwise. By Lemma 3.3(b),
$v_{p}\left(r_{1}/g_{1}^{2}\right)\leq 0$ $\leq v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)/p\right)$
$\leq v_{p}\left(\operatorname{rad}\left(\left(\operatorname{sf}\left(N_{\alpha}\right)/p\right)\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
for all $p\in I^{\prime}\cup I^{\prime\prime\prime}$ in the first case and for all $p\in I^{\prime\prime}\cup I^{\prime\prime\prime}$ in the second case. Hence with $f^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)/\max\left(d^{\prime}d^{\prime\prime\prime},d^{\prime\prime}d^{\prime\prime\prime}\right)$, we see that $f^{\prime}<\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ and that $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of
$4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right).$
(b) If $4b^{2}r_{1}/g_{1}^{2}$ is odd, then there is nothing to prove, as $4b^{2}r_{1}/g_{1}^{2}$ dividing $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ follows from $4b^{2}r_{1}/g_{1}^{2}$ being a divisor of $4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$. So we put $f=4b^{2}r_{1}/g_{1}^{2}$ and need only consider when $4b^{2}r_{1}/g_{1}^{2}$ is even.
@@
We can write $(r+si)/(1+i)=(r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
$-\left(\frac{2b^{2}r}{g_{1}^{2}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
Since $4b^{2}r/g_{1}^{2}$ is even, we know that $2b^{2}r/g_{1}^{2}\in\mathbb{Z}$.
Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}^{2}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$. Hence $2b^{2}r/g_{1}^{2}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ and so we take $f=2b^{2}r/g_{1}^{2}$ in this case.
@@
In this case, as well as cases (b-iii) and (b-iv), we will assume that $r_{1}=r_{1}^{\prime}$. The proof is identical using $r_{1}=r_{1}^{\prime\prime}$ instead.
@@
(b-ii-1) Suppose first that $v_{2}\left(r_{1}\right)$ is even.
In this case, $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$. If we have equality, then $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and so $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, as required. So we consider $v_{2}\left(4b^{2}\right)>v_{2}\left(r_{1}\right)$. Since both of these quantities are even, we must have $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)+2$. That is, $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$.
Since $s$ is even, we have
$v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)>2v_{2}\left(g_{1}\right).$
Since $r$ is odd, we also have $2v_{2}\left(g_{1}\right)=2v_{2}\left(r_{1}^{\prime}\right)$. Hence $v_{2}\left(r_{1}^{\prime\prime}\right)>v_{2}\left(r_{1}^{\prime}\right)$. As we stated near the start of this case (case (b-ii)), since $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$, we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
We saw in the proof of Lemma 3.2 that $v_{2}(y)\geq\min\left(v_{2}(b),v_{2}\left(r_{1}^{\prime}\right)/2\right)$. Since $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$, it follows that $v_{2}(y)\geq v_{2}\left(r_{1}^{\prime}\right)/2$. Hence
$v_{2}(2by)\geq 1+v_{2}\left(r_{1}\right)/2+v_{2}\left(r_{1}\right)/2=v_{2}\left(r_{1}\right)+1.$
Thus $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)\geq v_{2}\left(r_{1}\right)+1$, which implies that $v_{2}\left(r_{1}\right)>v_{2}\left(r_{1}\right)$. Therefore, the case when $v_{2}\left(r_{1}\right)$ is even and $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$ never occurs and we find that in case (b-ii-1), $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, as required.
(b-ii-2) Now suppose that $v_{2}\left(r_{1}\right)$ is odd.
Here $v_{2}\left(2b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$.
@@
Since $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)$, by the assumption that $s_{1}^{\prime}/g_{1}$ is odd, we have $v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)=v_{2}\left(g_{1}^{2}\right)$. Since $r$ is even, we also have $v_{2}\left(g_{1}^{2}\right)<2v_{2}\left(r_{1}\right)$, it follows that $v_{2}\left(r_{1}^{\prime\prime}\right)\neq v_{2}\left(r_{1}\right)$. Hence, by the comment at the start of case (b-ii), we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
(b-iii-1) If $v_{2}\left(r_{1}\right)$ is even, then $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}\right)$. Hence $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$ as required.
@@
Since $s$ is even, we have
$2\leq v_{2}\left(s^{2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)-v_{2}\left(2b^{2}r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime\prime}\right)-v_{2}\left(2b^{2}\right).$
So $v_{2}\left(r_{1}^{\prime\prime}\right)\geq 2v_{2}(b)+3$.
@@
From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4b^{2}N_{\varepsilon}$. So $atu+b^{2}du^{2}$ must be even. I.e., $atu$ and $b^{2}du^{2}$ have the same parity.
(b-iv-1) If $b^{2}du^{2}$ is odd, then $a$ and $t$ are also odd. Since $4N_{\varepsilon}=t^{2}-du^{2}$ is divisible by $4$ and $t$ and $u$ are both odd, it must be the case that $d\equiv 1\pmod{4}$.
If $d\equiv 1\pmod{8}$, then $4N_{\varepsilon}=t^{2}-du^{2}\equiv 0\pmod{8}$ and so $N_{\varepsilon}$ is even. Furthermore, since $abd$ is odd, we have $N_{\alpha}$ is even, so $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ is also even and the desired conclusion holds.
@@
(b-iv-2) We now consider $b^{2}du^{2}$ even. If $b$ is even, then
$v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)\geq 2>v_{2}(f)=1.$
So the desired conclusion holds. Therefore, we may assume that $b$ is odd and $du^{2}$ is even in what follows. We showed above that $uN_{\alpha}$ is even. We will show here that $N_{\varepsilon}$ is also even. These facts will suffice to show that $v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)\geq 1=v_{2}(f)$ and hence to prove part (b) in this case.
@@
Since $4N_{\varepsilon}=t^{2}-du^{2}\in\mathbb{Z}$ and $du^{2}$ is even, we have $t$ is even and hence $4|\left(du^{2}\right)$.
If $u$ is odd, then $4|d$. Also $N_{\alpha}$ is even, so $a$ is even. As a result, we find that $v_{2}(2atu),v_{2}\left(2b^{2}du^{2}\right)\geq 3$. So under our assumption that $N_{\varepsilon}$ is odd, from (3.4) we have $v_{2}\left(4y^{2}\right)=2$. I.e., $y$ is odd.
@@
In both cases, we find that $y$ is even. This contradicts what we obtained ($y$ is odd) under the assumption that $N_{\varepsilon}$ is odd. Hence we know that in the case when $b^{2}du^{2}$ is even, $N_{\varepsilon}$ is even and the assertion in part (b) of the lemma holds.
@@
(c-i) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=p$, for a prime, $p$. Then we have $f^{\prime}=1$ in part (a) and so $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$. Hence the result holds if $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
From Lemma 3.4, $r$ is even unless $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$. So, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$. With $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$.
(c-ii) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=2p$, where $p$ is an odd prime. From Lemma 3.4, we see that $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$ are both even. Hence, from Lemma 3.3(c), we have
$\max\left(v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{2}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)<0\leq v_{2}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
Similar to the end of the proof of part (a), we let $r_{1}=r_{1}^{\prime}$ if $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{2}\right)$ and $r_{1}=r_{1}^{\prime\prime}$ otherwise. From Lemma 3.3(b), we find that
$v_{p}\left(r_{1}/g_{1}^{2}\right)\leq 0\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
So with $f=4b^{2}r_{1}/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$. ∎